### v1

# Complex potential energy surfaces: gradients with projected CAP technique

Abstract
The complex absorbing potential (CAP) technique is one of the commonly used Non-Hermitian quantum mechanics approaches for characterizing electronic resonances. CAP combined with various electronic structure methods has shown promising results in quantifying the energies and widths of electronic resonances in molecular systems. While CAP-based methods can be used to map complex potential energy surfaces for resonance states, efficient exploration of these surfaces, e.g. geometry optimization or dynamical simulations, require information on the nuclear gradient. Currently, the only nuclear gradients available for CAP-based methods are for Hartree-Fock and Equation-of-Motion Coupled-Cluster method with single and double excitations (J. Chem. Phys. 146, 031101 (2017)). Here we provide a general approach that relies on projected CAP formulation and extends gradients and non-adiabatic couplings formulations developed for bound-state electronic structure methods to resonances. The approach is not limited to a specific electron structure method and is generally applicable to any electronic structure methods, provided the information on the gradients and non-adiabatic couplings is available for bound states. Here, we focus on the State-Averaged Complete Active Space Self-Consistent Field (SA-CASSCF) and Multi-Reference Configurational Interaction with Single excitation (MR-CIS) as our methods of choice. We establish the accuracy of the developed gradients and report equilibrium geometries for several representative temporary anion species (, , and ).
1. Introduction
Metastable electronic states are key precursors in reactive electron-molecule scattering processes. They play a key role in the breakage of single and double strands in DNA upon interaction with slow electrons1, 2, 3, 4 and have also been proposed as intermediates for the formation of stable anions in interstellar medium5, 6, 7, 8. These temporarily states, or resonances, lie in the electron detachment/ionization continuum, and their lifetimes vary from femtoseconds to picoseconds9, 10, 11, 12, which is long enough to be studied experimentally. Electronic resonances formed upon electron attachment can decay via several decay pathways accessible to the molecule. For short-lived anionic states, when the molecule does not leave the Franck-Condon region, an outgoing electron is emitted via autoionization and the molecule can end up in a vibrationally- excited state of the neutral, the process known as a resonance vibrational excitation (RVE)13, 14, 15. For longer-lived metastable states or steep dissociative surfaces for a metastable state, the molecule can undergo fragmentation, as is the case for dissociative electron attachments16, 17, 18, 19. Although single-point energy calculations can provide useful information on the energies and lifetimes of metastable states, a more thorough characterization of the processes and chemical reactions that proceed through electronic resonances requires tackling nuclear dynamics in metastable electronic states and disentangling nuclear motion and electronic decays, which often occur in comparable time-scapes20, 21, 22.
As resonances belong to the continuous spectrum of the electronic Hamiltonian, they cannot be straightforwardly described with bound-state methods. Special techniques have to be used20, 23, 24, 25, 26, 27. Moreover, a resonance is not represented by single electronic states in the Hermitian formulation28, 29, 30. However, from the standpoint of nuclear dynamics simulations, it is convenient to operate with a single electronic state that represents a resonance both for model development and results interpretation. The solution is provided by Non-Hermitian Quantum Mechanics (NHQM) methods 31, 23, 20. In this case, the resonances are associated with a single square-integrable state of modified non-Hermitian electronic Hamiltonian with complex Siegert-Gamow eigenvalue32, 33:
where and are geometry-dependent (R) resonance energy and width, respectively, forming complex potential energy surfaces (CPES).
Describing resonance with a single eigen-state is particularly attractive for mixed quantum/ classical dynamics models. In this case, chemical processes can be modeled by simulating nuclear motion classically34, 35 and if nuclear quantum effects are ignored36, the force vector can be taken as negative of the gradient of the real part of the complex energy, (R). The irreversible electronic decay into the continuum governed by the resonance width, (R). In fact, the first attempts have been made to use NHQM electronic structure methods in on-the-fly BO or surface hopping dynamics on complex potential energy surfaces35, 34. Apart from evaluating accurate complex energies, these methods necessitate evaluation of forces on complex surfaces and till now, analytic gradients have only been available for Hartree Fock (CAP-HF) and equation of motion coupled-cluster method for electron attachment with single and double substitution (CAP-EOM-CCSD) levels of theory37.
Here we present a general approach for computing analytic nuclear gradients for CPESs that uses analytic gradients and non-adiabatic couplings for discretized continuum states obtained from Hermitian electronic structure calculations as input. The approach relies on the projected complex absorbing potential (pCAP) method38, 39, 40, 41, which extends hermitian quantum mechanics methods to the NH domain. The pCAP method has been previously shown to produce accurate estimates of the resonance energies and widths for shape resonances in molecular systems38, 40, 41, 39. pCAP with careful choice of electronic structure methods has also been demonstrated to yield smooth CPESs42, 39, 43. Moreover, the quality of the pCAP CPESs was directly tested by comparing the computed cross-sections for resonant vibrational excitation (RVE) upon electron impact on N2 with the experimental ones: the cross-sections simulated with the wave-packed dynamics on the pCAP CPESs are in very good agreement with the experiment43.
The proposed strategy can be implemented for any electronic structure method, for which nuclear gradients and non-adiabatic couplings are available. Here, for the proof-of-concept implementation and for testing the accuracy of the resulting nuclear gradients we focus on two electronic structure methods of moderate computational cost: state-averaged complete active-space self-consistent field (SA-CASSCF) and multi-reference configuration interaction with single excitations (MR-CIS). While these methods, especially SA-CASSCF, might not sufficiently account for electron correlation, and, thus, might not yield accurate complex potential surfaces, the presented results show promise for CPESs nuclear gradient evaluations with more advanced quantum chemistry methods, e.g. multistate complete active space perturbation theory (MS-CASPT2)44, 45, 46, 47.
The structure of the article is as follows. Sec. 2. introduces pCAP method (Sec. 2.1.), describes the proposed model for analytic gradients evaluation (Sec. 2.2.), and summarizes the state tracking approach employed for geometry optimizations (Sec. 2.3.). Sec. 3. discusses system-specific details of the computational setup. Finally, computed analytic gradients are validated by comparison to the numerical ones (Sec. 4.1.) and by analyzing optimized geometries of representative metastable anions (Sec. 4.2.).
2. Theory
2.1. Projected CAP
CAP approach belongs to a group of NHQM methods, which describe an electronic resonance as a single eigen-state of a modified non-Hermitian Hamiltonian31, 23, 20. In CAP method, the Hermitian electronic Hamiltonian () is augmented by an imaginary absorbing potential ()48:
where is the CAP strength parameter, and is a one-particle operator specifying the functional form of CAP49, 50, 51. A resonance then appears as a single square-integrable eigen-state of with a complex eigenvalue (Eq. 1). Introduction of CAP absorbs the outgoing tail of the metastable wave function in the asymptotic limit and renders the wave-function square-integrable48. The most common choices are the cuboid 49 and Voronoi CAP50, 51. Here, we have used the cuboid CAP specified as follows:
where specifies the direction (), is the CAP onset in the corresponding direction, and is the value of the coordinate in a given direction. is the center of mass (COM) coordinate in direction for the molecular system defined as . and represents coordinate and mass of nucleus respectively. We use a center of mass oriented grid instead of a center of nuclear charges one37, which is merely a choice of convenience. As follows from Eq. 2, the , and consequently its eigen-values, depends on the strength parameter . In a complete basis set limit, the exact eigenvalues are recovered in the limit 48. For a finite basis, which is the case for practical calculations, one extracts the resonance position and width using an optimal value of , , obtained by analyzing stationary points of -trajectory, a series of complex energies computed for different values of . In this work, the was located using minimal logarithmic velocity criterion applied to uncorrected complex energies: 48.
By using CAP-augment Hamiltonian, , instead of regular Hermitian Hamiltonian one can extend any quantum chemistry method originally formulated for bound electronic states to resonances. Indeed, CAP approach has been used together with multiple single- and multi- reference electronic structure methods, including symmetry adapted cluster–configuration inter- action (SAC-CI)38, multi-reference perturbation theory (MRPT2)52, 53, equation of motion coupled-cluster (EOM-CC)54, 55, 56, 57, algebraic diagrammatic construction (ADC)58, 59, and multi-reference configuration interaction (MR-CI)60, 61, 62, 63. CAP approach can be integrated with an electronic structure method in two ways. First, one can operate with CAP-augmented Hamiltonian explicitly, which on practical level means that one has to modify the appropriate solvers and software to operate with complex Hamiltonians. Another consequence of explicit use of is the necessity of performing multiple electronic structure calculations to characterize a single resonance’s energy and width57, 64, 55, 42. This approach has been, for example, used to implement CAP-EOM-CCSD methods55. An alternative solution is provided by the projected CAP scheme, originally proposed by Ehara and Sommerfeld62. In this case, one starts with a regular Hermitian calculation employing Hermitian electronic Hamiltonian, (Eq. 2), and generates a basis of discretized continuum states, or zero-order (ZO) basis (). ZO states are, therefore, eigen-states of the parent Hermitian approximated using a given model chemistry. The is then projected on the ZO basis, resulting matrix
,where and are matrices of electronic Hamiltonian and in ZO basis. The former is most commonly diagonal (or can be made such, e.g. in the case of some flavors of multi-state perturbation theory). can be evaluated using one-particle reduced density and transition density matrices in atomic orbital (AO) basis () for ZO states as follows:
,where is CAP matrix calculated in AO basis.
Diagonalization of yields resonance as an eigen-state and resonance energy and width as real and imaginary parts of the corresponding complex eigenvalue62, 39. It has been shown that it is often enough to include only moderate number of states in ZO basis (10-30) to get a converged value of the resonance energy and width39. This approach has been used to implement CAP-based configuration interaction (CI)62, SAC-CI38, EOM-CCSD39, extended multiconfigurational quasidegenerate perturbation theory(XMCQDPT2)52, extended multistate complete active space perturbation theory (XMS-CASPT2)53 MR-CI41, and ADC65, 66.
2.2. Analytic nuclear gradients for pCAP electronic structure methods
The derivations below assume the Hermitian form of Hamiltonian such that has a complex-symmetric form. Yet, similar framework can be applied, for example, to non-Hermitian similarity-transformed Hamiltonian in EOM-CC methods by exploiting biorthogonal left and right eigen-states spaces. To set up the derivations, we first introduce a matrix, which is obtained by diagonalizing complex-symmetric by a complex orthogonal transformation matrix , a matrix of eigenvectors:
One of the diagonal matrix elements of is, therefore, the complex eigenvalue associated with the resonances. Owing to the complex-symmetric structure of left eigenvectors are transposed of the right ones, and, therefore, C matrix is orthogonal: . Equation (7) specifies transformation upon the basis change rotation from ZO to the diagonal basis, the basis of eigen-states of .
The nuclear gradient of the complex energy associated with the resonance can be, therefore, obtained from the gradient matrix written as follows:
With the two latter terms canceling out (see Sec. 2 of Supporting Information), the final expression relating the gradient matrix in diagonal and ZO basis has the following form:
Using Eq. (5) one can further rewrite the gradient matrix , referred to from here on as , by separating the terms originating from electronic Hamiltonian in ZO basis () and from CAP:
in Eq. 10 is the gradient matrix in ZO basis (), which has the following form:
where the diagonal elements and the off-diagonal elements are respectively the individual electronic state gradients and non-adiabatic couplings between the ZO states (), eigen-states of the Hermitian electronic Hamiltonian . Importantly, all of the matrix elements of the gradient matrix in ZO basis, , are obtained from conventional Hermitian electronic structure calculations in the absence of CAP and can be computed by standard quantum chemistry packages without any code modification.
The second term in the gradient matrix expression in diagonal basis, (Eq. 10), is the CAP contribution, which can be calculated numerically using a stand-alone software. A more detailed form of the Eq. (10) yields the following final form of the state’s gradient in the diagonal basis for n atom in direction:
where and are correlated basis state indices.
The quantity from the last term in equation (12) can be expressed similarly to eqn. (6) using the trace relation of one particle reduce density matrix in AO basis (). We further approximate that the is impervious to change in nuclear coordinate , leading to,
A closer inspection at the CAP gradient contribution in eqn. (13) reveals that it can be expanded into three contributions using eqn. (3),
where and are AO indices and is the atomic mass of nucleus.
One arrives to the first term by taking derivative of Wα (presented in eqn. (4)) with respect to the center of mass movement in direction. The analytical form of the looks as follows,
In equation (14), the last two parts correspond to contributions originating from the derivative of Gaussian basis function with respect to similar to those derived in Ref 37. Lastly, we keep the CAP onsets fixed in our geometry optimizations, hence, CAP derivative contributions with respect to are ignored37.
2.3. State tracking
Geometry optimization using a method operating with multiple electronic states requires a reliable state tracking approach to avoid root switching. In this work we used two state tracking approaches. The first is based on evaluating of the wave-functions overlap for two consecutive steps. Second method uses attachment and detachment densities for state tracking67, 68.
The overlap matrix is originally computed in the zero-order basis (). We used two approaches to calculate the overlap matrix. In first, the overlap was calculated by an external software WF-OVERLAP69 using information on molecular orbitals and the ZO states’ CI expansions as input. In the second, we used the built-in functionality of the OpenMolcas RASSI module70. The overlap matrix is then transformed to the diagonal basis () with the rotation matrices for the corresponding steps:
The optimization procedure follows the state that has the largest overlap with the state of interest from the previous step.
3. Computational details
Geometry optimizations were carried out using geomeTRIC71 and optking72 optimizers with the energy and gradients provided to the optimizer on-the-fly. For the systems with the symmetry point group higher symmetry than , the optimization was performed with optking software72 as symmetry is not implemented in geomeTRIC71. CAP strength parameter was kept fixed during optimization at value, optimal obtained for the equilibrium geometries of the neutral molecule. The box size was kept constant throughout the optimization. While this constraint can introduce some artifacts, it can be easily avoided by either using the Voronoi CAP50, 51 or adjusting the box CAP parameters along the optimization path. Exploring the effects of CAP onto the optimized equilibrium geometries of resonances is the subject of the future work. The complex energies evaluated with several considered model chemistries were computed with OpenCAP software interfaced73 to OpenMolcas70 and COLUMBUS74 electronic structure packages. Zero-order gradients and non-adiabatic couplings for SA-CASSCF and MR-CIS calculations were computed with OpenMolcas70 and COLUMBUS74 packages, respectively.
Our main choice of parent valence basis set is cc-pVTZ basis that has been further augmented by a subset of diffuse basis functions. Two types of diffuse subsets have been used. In the first, a single ghost atom (denoted as X) was placed in the center of mass of the molecule for each nuclear configuration (for every optimization step). In the second, the parent basis on the heavy atoms were augmented with several even-tempered diffuse basis functions. The specific basis sets used for each of the considered electronic resonances are given in the corresponding section below.
The computational details specific to each studied system are summarized below.
3.1. anion: resonance
Equilibrium geometry of the anion in the resonance was obtained using two methods: State-averaged CASSCF (SA-CASSCF) and MR-CIS. The geometry optimization was performed within symmetry point group. State-averaging over 10 states was employed for SA-CASSCF: one state representing the neutral and 9 anionic states of symmetry to represent the continuum. The active space chosen for this calculation was 5 electrons in 13 orbitals (1, 1, 5, 5, 1), where orbitals are represented by degenerate pairs of and orbitals.
MR-CIS calculations were carried out with smaller active spaces: (5e, 9o) and (7e,10o). Active space orbitals included for these two choices are (1, 1, 3, 3, 1) and (1, 1, 1, 3, 3, 1) respectively. State-averaging over six states was used in both cases with one state and five anionic states representing the neutral and discretized continuum, respectively.
MR-CIS calculations for the neutral and anionic states were performed using the same SA-CASSCF orbitals to get a balanced description of the neutral and anionic resonance states. No orbitals were frozen in the calculations43. The starting N-N bond length for these three geometry optimizations is taken to be 1.095 Å , corresponding to CASSCF(2e,2o) optimization of the neutral molecule with cc-pVTZ basis set.
To compare the equilibrium bond length and resonance energetic parameters with the previously reported data41, 43 and precomputed MR-CIS curves41, we also performed a SA-5-MR-CIS calculation with (5e, 8o) active space using the recipe prescribed by Thodika and Matsika41. All the electronic structure calculations, including ZO energies75, densities, gradients and non-adiabatic couplings76, 77 evaluations were performed with COLUMBUS electronic structure package74.
3.2. anion: resonance
SA-CASSCF geometry optimization of formaldehyde anion was performed with no symmetry constraints. Several state-averaging schemes have been used. In all cases, the state-averaging was performed over anionic states. The converged SA-CASSCF orbital were then used to perform CASCI calculation for the ground state of the neutral to get the reference energies. This choice of state-averaging scheme can introduce imbalance in description of the neutral and anion, which in turn can affect the computed resonance position. Yet, one can expect that not including the neutral in the state-averaging does not lead to detrimental description of the complex potential energy surface of the anionic resonance, and therefore, it’s equilibrium geometry. As the primary focus of this work is on nuclear gradients for the complex potential energy surfaces, this choice of methodology does not affect the main conclusions of the work. Two diffuse basis sets centered on the ghost atoms were used. The first basis set consisted of four -type basis functions (4p). The second basis included two -, five - and two - type even-tempered basis functions. In both cases, the initial exponents were chosen based on the most diffuse exponent in the parent basis. The scaling factor of two was used to generate the exponents of the even-tempered functions.<– not true always, we have used a combination of 1.5 and 2.0. The scaled exponents are available in the Supporting Information
MR-CIS geometry optimization was carried out using the symmetry point group. (3e, 8o) CASSCF active space consisted of 1 () and 7() orbitals with 2 electrons in and an extra electron in (). The state averaging over 7 () anionic states was used. These CASSCF orbitals were used to perform subsequent anionic() and neutral () MR-CIS calculations.
3.3. Formic acid anion: resonance
Calculations for formic acid anion were performed with SA-CASSCF and MR-CIS. To assess the the basis set effects on the optimized geometry several different diffuse basis subsets were used: 3p, 4p, 5p, and 2s5p2d bases for SA-CASSCF. The exponents are available in the Supporting Information.
For MR-CIS calculation, cc-pVDZ and cc-pVTZ basis sets were used with a ghost atom in the center of mass carrying [3p] and [4p] diffuse orbitals. (3e, 12o) active space was used for both basis sets. State-averaging over seven and eight states were used for cc-pVTZ+[3p] and cc-pVDZ+[4p] respectively.
To explore the effects of interaction with the continuum on the geometry of the anion, we have also performed a reference geometry optimization for valence localized state of formic acid, which was achieved by using the basis without any diffuse basis functions. The active space chosen for the latter calculation was 3 electrons over 4 orbitals and the valence root of SA-CASSCF and subsequent MR-CIS were optimized.
3.4. anion: resonance
CASSCF level geometry optimization was carried out with no symmetry constraints. cc-pVTZ basis augmented with three extra diffuse -type basis functions placed on the carbon atoms was used. A preliminary CASSCF/cc-pVTZ optimized geometry of the neutral was chosen as a starting point for the geometry optimization of the anion.
MR-CIS geometry optimization was performed using the symmetry point group. Same basis set as discussed above was employed. The starting geometry was the same as for the SA-CASSCF geometry optimization. (3e, 10o) active space included one and nine orbitals. State-averaging over eight states of symmetry was used. The reference MR-CIS energy of the neutral was evaluated using the converged CASSCF orbitals.
A smaller cc-pVDZ basis with three extra diffuse -type orbitals placed on the carbon atoms was also used to assess the basis set effect on the equilibrium geometry of the anion. The (3e,7o) SA-CASSCF active space in this case included 1 and 6 orbitals. State averaging over six states of symmetry was used. The same approach to that used for a larger basis was taken to evaluate the MR-CIS reference energy of the neutral.
Similarly to formic acid anion, to explore the effects of the interaction with the continuum on the equilibrium geometry of the resonance of C2H, CASSCF(2e,2o) geometry optimization for the valence localized anion was performed using cc-pVTZ basis.
4. Results
Below we discuss the performance and accuracy of the proposed form of the analytic gradients for the complex potential energy surfaces for pCAP-based methods. We first focus on comparison between analytic and numerical gradients. We then discuss the optimized geometries of representative metastable anions.
4.1. Validation against the numerical gradients
To test the accuracy of the proposed analytic gradients, we have performed the comparison between analytic and numerical gradients. Note that, as approximations have been introduced, i.e. the variation in the rotation matrix with the change in the nuclear coordinates for evaluating the CAP contribution, we do not anticipate the exact agreement. For numerical calculation of the gradient, we choose a R =0.0001Å step in each direction for all the atoms and calculate the gradient using two-point finite difference. Fig. 1 shows the computed numerical and analytic gradients for MR-CIS (N) and SA-CASSCF (formic acid anion). We see a good agreement between the two sets of data, which points to a negligible contributions of the approximations made. Moreover, comparing the gradient terms originating from CAP for CAP-EOM-EA-CCSD37 and pCAP-EOM-EA-CCSD points to only minor differences, which also supports the proposed form of the gradients (see Supporting Information).
4.2. Equilibrium geometries of metastable anions
We have considered four representative shape resonances in molecular systems to explore the performance of the pCAP methods and the corresponding analytic gradients for predicting equilibrium geometries of metastable anions. Specifically, we have considered the shape resonances in the anions of dinitrogen (), formaldehyde (), formic acid () and ethene (). The resulting equilibrium geometries and relevant natural orbitals are shown in Figure 2. Below we discuss each of the resonances separately.
singlelinecheck=off aMR-CIS/SA-6-CASSCF(7e,10o), cc-pVTZ+[2s5p2d] (D2h symmetry); bMR-CIS/SA-7-CASSCF(3e,8o)/cc-pVDZ+[3p] (Cs symmetry); cSA-15-CASSCF(3e,23o)/cc-pVTZ+X[5p] (C1 symmetry); dMR-CIS/SA-6-CASSCF(3e,7o)/cc-pVTZ+[3p] ( symmetry).
4.2.1. shape resonance
Optimized bond lengths and energetic parameters for the equilibrium geometries of the neutral and of the shape resonance are listed in Table 1. Population of the orbital leads to the effective bond order decrease and bond elongation. Indeed, all method and basis set combinations correctly capture this behavior: the bond length increases by 0.06 - 0.10 Å depending on the method. As mentioned earlier, the lack of dynamical electron correlation in SA-10-CASSCF(5e, 13o) leads to inaccurate description of the resonance energetic parameters. The resonance energy of 4.4 eV at the equilibrium geometry of the neutral is more than 1 eV away from the experimental value78 and from more accurate theoretical estimates62, 64, 39, 41. In addition, SA-10-CASSCF(5e, 13o) significantly overestimates the width of the resonance (Table 1). Taking into account dynamic electron correlation brings the resonance down in energy. MR-CIS/SA-6-CASSCF(5e, 9o) yields resonance energy and width at the equilibrium of the neutral of 3.2 and 0.54 eV, respectively, which are much closer to the experimental estimate78. The equilibrium bond length of the neutral is identical to CASSCF results. The bond length for the optimized geometry of the anion is slightly greater than the corresponding SA-10-CASSCF(5e, 13o) value (by 0.02Å). We have also explored the effects of extending the active space by including additional occupied molecular orbital in the active space yielding MR-CIS/SA-6-(7e, 10o) model. The resulting changes in resonance energy, width, and equilibrium bond length are minor (Table 1). To compare the optimized geometries to previously reported data obtained with similar methods, we have performed MR-CIS/SA-10-CASSCF(5e,8o) calculations. One of the difference in the data reported in Ref.41 and here is the number of ZO states used for pCAP: five states used here vs. ten states in Ref. 41. Thodika and Matsika have reported a shallow potential energy minimum close to 1.15Å by scanning along the N–N distance with 0.05Å step. A cubic interpolation of their potential energy curve data yields a minima at 1.17Å . Using a similar MR-CIS/SA-10-CASSCF(5e, 8o) approach, we locate the minimum at 1.18Å . The energies and the widths are very close to those reported in Ref. 41 for both equilibrium geometries of the neutral and of the anion.
Geometry relaxation in the anionic state brings the two states closer in energy by 0.5 - 1 eV depending on the methods, which is also reflected in the consistent decrease in the resonance width (Table 1).
| Methods/basis | N2 | N | ||||
| E () | () | E () | () | |||
| SA-10-CASSCF(5e, 13o) | 1.095 | 4.448 | 1.622 | 1.155 | 3.982 | 1.070 |
| cc-pVTZ+[2s5p2d] | ||||||
| MR-CIS/SA-6-CASSCF(5e, 9o) | 1.095 | 3.166 | 0.544 | 1.176 | 2.418 | 0.346 |
| cc-pVTZ+[2s5p2d] | ||||||
| MR-CIS/SA-6-CASSCF(7e, 10o) | 1.095 | 3.516 | 0.631 | 1.193 | 2.538 | 0.376 |
| cc-pVTZ+[2s5p2d] | ||||||
| MR-CISa/SA-10-CASSCF(5e,8o) | 1.100b | 3.033 | 0.503 | 1.180 | 2.226 | 0.329 |
| aug-cc-pVTZ+X[3s3p3d] | ||||||
4.2.2. Formaldehyde anion: shape resonance
Table. 2 lists the geometric parameters for the optimized geometries of the formaldehyde in the ground state of the neutral and in the shape resonance state. As anticipated, populating the orbital leads to elongation of the C=O double bond: the bond length increases by 0.09-0.10Å depending on the electronic structure method. The H-C-O angles and C–H bond lengths do not change considerably. Overall, the geometries obtained with both SA-CASSCF and MR-CIS are quite similar.
| Methods/basis | CH2O | CH2O- | ||||
| R(C=O) | R(C–H) | (H–C–O) | R(C=O) | R(C–H) | (H–C–O) | |
| SA-15-CASSCF(3e, 16o)/ | 1.178 | 1.093 | 122 | 1.278 | 1.110 | 120 |
| cc-pVTZ+X[4p] | ||||||
| SA-19-CASSCF(3e, 20o)/ | 1.178 | 1.093 | 122 | 1.268 | 1.098 | 121 |
| cc-pVTZ+X[2s5p2d] | ||||||
| MRCIS/SA-7-CASSCF(3e, 8o)/ | 1.182 | 1.101 | 122 | 1.282 | 1.102 | 122 |
| cc-pVDZ+[3p] (Cs symmetry) | ||||||
The resonance parameters for the equilibrium geometry of the neutral and of the shape resonance are listed in the Table 3. Similarly to what has been observed for the N2, SA-CASSCF significantly overestimates resonance energy at the equilibrium geometry of the neutral placing the state at 2.4- 2.6 eV above the neutral. Including of the dynamic electron correlation effects brings the resonance energy down to 0.9 eV, which is in much better agreement with earlier more accurate theoretical estimates38, 79, 37, 80 and with the experimental reference value of 0.8681. Interestingly, the widths obtained with the three methods rather close (variation within 0.2 eV). Notably, while extension of the diffuse basis set from [4p] to [2s5p2d] leads to the shift in resonance energy of 0.2 eV and to the increase in the width of 0.08 eV, the effects on the optimized geometry are negligible (Table 2). Same trends for resonance energetic parameters, i.e. SA-CASSCF overestimating the resonance position and MR-CIS bringing the energy down, are observed for other molecular shape resonances, and, therefore, below we only focus on the equilibrium geometries rather than energetic parameters, which are reported in Supporting Information.
| Methods/basis | ||||
| SA-15-CASSCF(3e, 16o)/ | 2.426 | 0.338 | 1.281 | 0.232 |
| cc-pVTZ+X[4p] | ||||
| SA-19-CASSCF(3e, 20o)/ | 2.633 | 0.416 | 1.086 | 0.687 |
| cc-pVTZ+X[2s5p2d] | ||||
| MRCIS/SA-7-CASSCF(3e, 8o)/ | 0.931 | 0.211 | 0.168 | 0.101 |
| cc-pVDZ+[3p] | ||||
4.2.3. Formic acid anion: shape resonance
Formic acid undergoes significant structural changes in the resonance relative to the neutral37, 82. In addition to the single C–O, double C=O bonds elongation geometry relaxation is accompanied by the out-of-plane distortion37, 82. Table 4 reports the equilibrium values of the relevant structural parameters for the neutral and anionic states of formic acid.
All methods consistently reproduce the increase in the C=O and C–O bond lengths of 0.08 - 0.09Å and 0.10 - 0.14Å respectively. Most variation between the methods is observed in the predicted magnitude of the H-C-O-H dihedral angle. pCAP-based SA-CASSCF methods yield the angle of 116-118 ∘, while pCAP-based MR-CIS yields the value of 134∘. Owing to these notable structural changes, we have chosen formic acid anion as a model system to explore the effects of the size of the diffuse basis subset on the resulting structural parameters. As follows from Table 4 addition of extra diffuse basis functions beyond 3-type functions on the ghost atom does not lead to significant changes in the optimized geometry configurations of the anion. Importantly, the SA-3-CASSCF(3e, 4o) geometry optimization approximating the resonance by a localized valence state (penultimate row in the Table 4) yields bond lengths and the dihedral angle that are rather different from the comparable pCAP SA-CASSCF calculations. The localized simulation overestimated the deviation from the planarity with the resulting H-C-O-H 134∘ in comparison to 116-118∘ for pCAP-based calculations. Similar trend is observed for MR-CIS method, where geometry optimization of a localized resonance yields a structure with a stronger out-of-plane relative to the pCAP-based calculation: 140∘ vs 134∘, respectively. These discrepancies in the optimized structural parameters of the resonance and approximate localized state point to the necessity of the proper account of the metastable character of the state.
These set of observations establish several importance of pCAP method based optimizations. Firstly, representing resonance state by bound state calculation does not provide the same description of the resonance. The geometrical parameters compared in the Table 4 shows that even though it may be possible to represent the resonance state by valence bound state, it does not necessarily guarantee valid results. Secondly, the convergence with respect to the basis sets size provides the glimpse that just adding 3 diffuse orbitals in the ghost atom is reasonably good enough to describe the the discretized continuum.
| Methods/basis | HCOOH | HCOOH- | ||||
| R(C=O) | R(C–OH) | (H–C–O-H) | R(C=O) | R(C–OH) | (H–O-C–H) | |
| SA-16-CASSCF(3e, 17o)/ | 1.178 | 1.319 | 180 | 1.264 | 1.434 | 117 |
| cc-pVTZ+X[3p] | ||||||
| SA-15-CASSCF(3e, 23o)/ | 1.178 | 1.319 | 180 | 1.261 | 1.428 | 118 |
| cc-pVTZ+X[4p] | ||||||
| SA-15-CASSCF(3e, 23o)/ | 1.178 | 1.319 | 180 | 1.267 | 1.426 | 118 |
| cc-pVTZ+X[5p] | ||||||
| SA-16-CASSCF(3e, 17o)/ | 1.178 | 1.319 | 180 | 1.265 | 1.430 | 116 |
| cc-pVTZ+X[2s5p2d] | ||||||
| MRCIS/SA-7-CASSCF(3e, 12o)/ | 1.178 | 1.319 | 180 | 1.269 | 1.439 | 132 |
| cc-pVTZ+X[3p] | ||||||
| MRCIS/SA-8-CASSCF(3e, 12o)/ | 1.180 | 1.322 | 180 | 1.267 | 1.465 | 134 |
| cc-pVDZ+X[4p] | ||||||
| SA-3-CASSCF(3e, 4o)/ | 1.178 | 1.319 | 180 | 1.248 | 1.420 | 134 |
| cc-pVTZ (valence-, w/o CAP) | ||||||
| MR-CIS/SA-3-CASSCF(3e, 4o)/ | 1.178 | 1.319 | 180 | 1.265 | 1.462 | 140 |
| cc-pVTZ (valence-, w/o CAP) | ||||||
4.2.4. : shape resonance
Finally, we present the results of the geometry optimization of the ethylene anion in the resonance state. SA-CASSCF as well as the two MR-CIS calculations yield very similar results. Both sets of calculations reproduce bond extension in the state of 0.08-0.10 Å. The dihedral angle is slightly higher for SA-15-CASSCF (37∘) in comparison to the MR-CIS (30∘). of the Geometry optimization with and without symmetry has almost similar geometry change for ethylene temporary anion. The distortion from planarity is also observed and is even more pronounced in the CAP-free calculations (42∘). Benda and Jagau37 has reported a dihedral angle change of 27∘ for a CAP-EOM-EA-CCSD based optimization with aug-cc-pVDZ+3s3p(A) basis, which is quite close to MR-CIS results reported here.
| Methods/basis | ||||
| R(C=C) | (H–C–C-H) | R(C=C) | (H–C–C-H) | |
| SA-15-CASSCF(3e, 20o)/ | 1.330 | 0 | 1.429 | 37 |
| cc-pVTZ+[3p] | ||||
| MR-CIS/SA-8-CASSCF(3e, 9o)/ | 1.341 | 0 | 1.421 | 30 |
| cc-pVTZ+[3p] ( symmetry) | ||||
| MR-CIS/SA-6-CASSCF(3e, 7o)/ | 1.341 | 0 | 1.426 | 31 |
| cc-pVDZ+[3p] ( symmetry) | ||||
| CASSCF(3e, 2o)/ | 1.330 | 0 | 1.437 | 42 |
| cc-pVTZ (valence-, w/o CAP) | ||||
Overall, proposed form of the analytic gradients allows for geometry optimization on complex potential energy surfaces with a variety of electronic structure methods. Here we demonstrate the use of CASSCF and MR-CIS approaches. The reported optimized geometries of the metastable anions point to differences in the approximate localized treatment of the resonance state and CAP-based approaches, pointing to the necessity of appropriate treatment of interaction with the continuum.
4.3. Conclusions
The described formulation of the analytic nuclear gradients allows one to extend the methods and software for calculating gradients and non-adiabatic couplings for bound electronic states to resonances. The computed analytic gradients are in excellent agreement with their numerical counterparts, despite some terms have been neglected. In this work, we have presented implementation of analytic gradients for two projected CAP based electronic structure methods: CASSCF and MR-CIS. While CASSCF, as can be expected, yields poor energetic parameters owing to the lack of account for dynamic electron correlation, optimized geometries capture the qualitative changes observed with a more reliable MR-CIS method. The described scheme also produces non-adiabatic couplings as a by-product, i.e. as off-diagonal matrix elements of the gradient matrix in the diagonal basis. The described developments pave the way for on-the-fly Born-Oppenheimer and non-adiabatic dynamics involving metastable states.
5. Supporting Information
The Supporting Information contains information on the details of the numerical analysis of the presented pCAP gradient and on the comparison with other methods, presents energetic parameters of all discussed shape resonances, and lists diffuse basis sets used for calculation. The OpenCAP software with CAP gradient implementation is available in https://github.com/SoubhikM/opencap/ with geometry optimization interfaces with COLUMBUS and OpenMolcas in https://github.com/SoubhikM/OPENCAP-MD/.
6. Acknowledgments
The authors thank Dr. Mushir Thodika and Professor Spiridoula Matsika for providing their MR-CIS potential energy curve data for . Simulations were performed on Boston University’s Shared Computing Cluster (SCC).

### v2

@@
 In both cases, the initial exponents were chosen based on the most diffuse exponent in the parent basis.
-The scaling factor of two was used to generate the exponents of the even-tempered functions.<– not true always, we have used a combination of 1.5 and 2.0.
 The scaled exponents are available in the Supporting Information
@@
 | cc-pVDZ+[3p] (Cs symmetry) | ||||||
-The resonance parameters for the equilibrium geometry of the neutral and of the shape resonance are listed in the Table 3.
+The resonance parameters for the equilibrium geometry of the neutral and of the shape resonance are listed in the Table S3.
 Similarly to what has been observed for the N2, SA-CASSCF significantly overestimates resonance energy at the equilibrium geometry of the neutral placing the state at 2.4- 2.6 eV above the neutral.
@@
 Simulations were performed on Boston University’s Shared Computing Cluster (SCC).
+Supporting Information
+Complex potential energy surfaces: gradients with projected CAP technique
+Soubhik Mondal and Ksenia B. Bravaya
+Department of Chemistry, Boston University, Boston, Massachusetts, 02215, USA
+S1. Comparison of CAP contributions to the gradient: CAP-EOM-EA-CCSD vs pCAP-EOM-EA-CCSD
+| CAP-EOM-EA-CCSD | pCAP-EOM-EA-CCSD | |||||
+| , : (2.605, 0.542) eV | , : (2.605, 0.546) eV | |||||
+| Atom | X | Y | Z | X | Y | Z |
+| N1 | 0.00000000 | 0.00000000 | 0.0567769 | 0.00000000 | -0.00000000 | 0.05810603 |
+| N2 | 0.00000000 | 0.00000000 | -0.0567769 | 0.00000000 | -0.00000000 | -0.05810603 |
+In Table S1, we report CAP contributions to the gradient for a single point calculation of resonance in CAP-EOM-EA-CCSD1 and pCAP-EOM-EA-CCSD2 levels respectively.
+The similarities in magnitudes and signs of the CAP gradients from two different methods underlines that firstly, complex 1-RDMs from our method (S1) is similar to CAP-EOM-EA-CCSD where CAP is introduced in HF level and secondly, the evaluation of CAP derivative terms implementation on the grid is correct.
+We don’t expect the complex densities and CAP terms to be exactly same even though the complex energies are same because they are entirely different methods.
+In ‘diag’ basis, we evaluate the 1-particle reduce density matrices (1-RDM) as following,
+,where and ,where are the 1-TDM (transition density matrices) between state I and J in ‘diag’ and ZO basis respectively. If I and J are the same, they correspond to state density matrices.
+S2. Rotation vector contribution to the gradient
+We utilize the orthogonality relation of the rotation matrix : and the eigenvalue equation: , , where are the complex eigen-values. It follows that,
+S3. Tracking resonance state: using attachment and detachment densities
+The tracking algorithm we deploy when full wave-function overlap is not available, is through comparison of attachment and detachment densities3, 4 in subsequent steps, an approach which is also adopted in the Q-CHEM program5.
+Equation-wise, the overlap metric between state I and J from consecutive steps is evaluated as following4,
+where, and are attachment and detachment densities in ‘diag’ basis for old and new iterations respectively.
+In equation (S3), the norm of the matrix M is evaluated as with being eigenvalues of .
+Essentially, equation (S3) provides an estimate of how the density of a state changes between two iterations, if the densities are close, is 1.
+The attachment and detachment densities for state are calculated using prescriptions from reference 3 and the one particle difference density matrix ()3 is built with the reference state chosen as the neutral state evaluated in the same orthogonal basis to that of the correlated anionic states.
+The density contraction to ‘diag’ from ZO basis is done using equation (S1).
+S4. pCAP parameters for initial calculations at equilibrium geometries of the neutral & energetics
+| Methods/basis | ||||
+| SA-10-CASSCF(5e, 13o) | 2.77 | 2.77 | 4.88 | 0.00790 |
+| cc-pVTZ+[2s5p2d] | ||||
+| MR-CIS/SA-6-CASSCF(7e, 10o) | 2.77 | 2.77 | 4.88 | 0.00854 |
+| cc-pVTZ+[2s5p2d] | ||||
+| MR-CIS/SA-6-CASSCF(5e, 9o) | 2.77 | .2.77 | 4.88 | 0.00724 |
+| cc-pVTZ+[2s5p2d] | ||||
+| MR-CIS/SA-10-CASSCF(5e, 8o) | 2.76 | 2.76 | 4.88 | 0.00826 |
+| aug-cc-pVTZ+X[3s3p3d] |
+| Methods/basis | ||||||
+| SA-15-CASSCF(3e, 16o)/ | 5.98 | 2.426 | 0.338 | 0.01656 | 1.281 | 0.232 |
+| cc-pVTZ+X[4p] | 3.88 | |||||
+| 2.93 | ||||||
+| SA-19-CASSCF(3e, 20o)/ | 5.99 | 2.633 | 0.416 | 0.02888 | 1.086 | 0.687 |
+| cc-pVTZ+X[2s5p2d] | 3.89 | |||||
+| 2.94 | ||||||
+| MRCIS/SA-7-CASSCF(3e, 8o)/ | 6.05 | 0.931 | 0.211 | 0.00196 | 0.168 | 0.101 |
+| cc-pVDZ+[3p] (Cs symmetry) | 3.94 | |||||
+| 2.97 | ||||||
+| Methods/basis | ||||||
+| SA-15-CASSCF(3e, 20o)/ | 7.62 | 3.119 | 0.634 | 0.01772 | 2.350 | 0.364 |
+| cc-pVTZ+X[3p] | 4.20 | |||||
+| 6.02 | ||||||
+| MR-CIS/SA-6-CASSCF(3e, 7o)/ | 7.14 | 2.617 | 0.592 | 0.00465 | 2.079 | 0.348 |
+| cc-pVDZ+[3p] ( symmetry) | 3.50 | |||||
+| 4.69 | ||||||
+| Methods/basis | ||||||
+| SA-16-CASSCF(3e, 17o)/ | 9.45 | 3.889 | 0.199 | 0.00743 | 1.591 | 0.258 |
+| cc-pVTZ+X[2s5p2d] | 5.29 | |||||
+| 3.56 | ||||||
+| SA-16-CASSCF(3e, 17o)/ | 9.45 | 4.116 | 0.240 | 0.00864 | 1.580 | 1.125 |
+| cc-pVTZ+X[3p] | 5.28 | |||||
+| 3.52 | ||||||
+| SA-15-CASSCF(3e, 23o)/ | 9.54 | 3.739 | 0.065 | 0.04005 | 1.636 | 0.103 |
+| cc-pVTZ+X[4p] | 6.36 | |||||
+| 6.33 | ||||||
+| SA-15-CASSCF(3e, 23o)/ | 9.54 | 4.160 | 0.133 | 0.03878 | 1.674 | 0.496 |
+| cc-pVTZ+X[5p] | 5.29 | |||||
+| 3.56 | ||||||
+S5. Augmented diffuse basis
+The main parent basis used in our calculations is cc-pVTZ, unless mentioned this is the case for the following.
+S5.1.
+2s,2d scaling exponent of . 5p scaling exponent of .
+ΨNΨΨΨ0 ΨS 1 1.00 Ψ0.0893500 1.0000000 ΨS 1 1.00 Ψ0.0446750 1.0000000 ΨP 1 1.00 Ψ0.1150 1.0000000 ΨP 1 1.00 Ψ0.0766667 1.0000000 ΨP 1 1.00 Ψ0.0511111 1.0000000 ΨP 1 1.00 Ψ0.0340741 1.0000000 ΨP 1 1.00 Ψ0.0227160 1.0000000 ΨD 1 1.00 Ψ0.234500 1.000000 ΨD 1 1.00 Ψ0.117250 1.000000
+3s,3p.3d scaling exponent for for aug-cc-pVTZ.
+ΨXΨΨ0 ΨS 1 1.00 Ψ0.0288 1.00000000 ΨS 1 1.00 Ψ0.0144 1.00000000 ΨS 1 1.00 Ψ0.0072 1.00000000 ΨP 1 1.00 Ψ0.02455 1.0000000Ψ ΨP 1 1.00 Ψ0.012275 1.00000000 ΨP 1 1.00 Ψ0.0061375 1.00000000 ΨD 1 1.00 Ψ0.0755 1.000000 ΨD 1 1.00 Ψ0.03775 1.00000 ΨD 1 1.00 Ψ0.018875 1.00000
+S5.2.
+3p scaling exponent with cc-pVDZ.
+ΨCΨΨ0 ΨP 1 1.00 Ψ0.13765 1.0 ΨP 1 1.00 Ψ0.068825 1.0 ΨP 1 1.00 Ψ0.0344125 1.0
+4p scaling exponent
+ΨX 0 ΨP 1 1.00 Ψ0.0806 1.000000D+00 ΨP 1 1.00 Ψ0.05373 1.000000D+00 ΨP 1 1.00 Ψ0.03582 1.000000D+00 ΨP 1 1.00 Ψ0.02388Ψ1.00
+2s,2d scaling exponent of . 5p scaling exponent of .
+ΨX 0 ΨS 1 1.00 Ψ0.05135Ψ1.0 ΨS 1 1.00 Ψ0.025675 1.0000000 ΨP 1 1.00 Ψ0.0806 1.000000D+00 ΨP 1 1.00 Ψ0.05373 1.000000D+00 ΨP 1 1.00 Ψ0.03582 1.000000D+00 ΨP 1 1.00 Ψ0.02388Ψ1.00 ΨP 1 1.00 Ψ0.01592Ψ1.00 ΨD 1 1.00 Ψ0.159 1.0000000 ΨD 1 1.00 Ψ0.0795 1.0000000
+S5.3.
+3p scaling exponent with cc-pVDZ.
+ΨP 1 1.00 Ψ0.13765 1.0 ΨP 1 1.00 Ψ0.068825 1.0 ΨP 1 1.00 Ψ0.0344125 1.0
+3p scaling exponent .
+ΨCΨ0 ΨP 1 1.00 Ψ0.0806Ψ1. ΨP 1 1.00 Ψ0.05373333333333333Ψ1. ΨP 1 1.00 Ψ0.03582222222222222Ψ1.
+S5.4.
+4p/5p: scaling exponent of .
+X 0 P 1 1.00 0.06045 1.000000D+00 P 1 1.00 0.030225 1.000000D+00 P 1 1.00 0.0151125 1.000000D+00 P 1 1.00 0.00755625Ψ 1.000000D+00 P 1 1.00 0.003778125 1.000000D+00
+3p, scaling exponent .
+X 0 P 1 1.00 0.0806 1.000000D+00 P 1 1.00 0.05373 1.000000D+00 P 1 1.00 0.03582 1.000000D+00
+2s,2d scaling exponent is ). 5p scaling exponent is ..
+X 0 S 1 1.00 0.05135Ψ1.0 S 1 1.00 0.025675 1.0000000 P 1 1.00 0.0806 1.000000D+00 P 1 1.00 0.05373 1.000000D+00 P 1 1.00 0.03582 1.000000D+00 P 1 1.00 0.02388Ψ1.0 P 1 1.00 0.01592Ψ1.0 D 1 1.00 0.159 1.0000000 D 1 1.00 0.0795 1.0000000
+S6. Initial geometries
+S6.1. Initial geometry (in Å) of with cc-pVTZ basis
+with symmetry.
+N 0.00000000 0.00000000 0.54743386 N 0.00000000 0.00000000 -0.54743386
+S6.2. Initial geometry (in Å) of with cc-pVTZ basis
+with symmetry.
+C 0.539081 -0.009227 0.000000 O -0.636180 -0.086613 0.000000 H 1.181355 -0.893267 0.000000 H 1.056513 0.968385 0.000000
+S6.3. Initial geometry (in Å) of with cc-pVDZ basis
+with symmetry.
+C 0.535543212514 -0.008891235494 0.00000000 O -0.643690629882 -0.088924526013 0.00000000 H 1.185776060252 -0.897982505052 0.00000000 H 1.063138712502 0.975076216165 0.00000000
+S6.4. Initial geometry (in Å) of with cc-pVTZ basis
+with symmetry.
+C 0.150321 -0.378974 0.000000 H 0.213778 -1.461673 0.000000 O 1.082107 0.341490 0.000000 O -1.109733 0.010919 0.000000 H -1.143257 0.957261 0.000000
+S6.5. Initial geometry (in Å) of with cc-pVDZ basis
+with symmetry.
+C 0.151159 -0.375236 0.000000 H 0.213978 -1.480920 0.000000 O 1.079429 0.352747 0.000000 O -1.113650 0.010337 0.000000 H -1.137699 0.962095 0.000000
+S6.6. Initial geometry (in Å) of with cc-pVTZ basis
+with symmetry.
+C -0.0179750928 -0.1136932028 -0.0115175428 H -0.0496468803 0.0145973544 1.0544147734 H 0.9369204904 0.0146037045 -0.4862662659 C -1.1084247201 -0.4170070183 -0.7097827982 H -1.0767529326 -0.5452975755 -1.7757145852 H -2.0633203033 -0.5453028673 -0.2350335460
+S6.7. Initial distorted geometry (in Å) of with cc-pVTZ basis
+with symmetry.
+ΨC 0.670519 -0.000015 0.000000 ΨC -0.670519 0.000015 0.000000 ΨH 1.238290 0.000271 0.922918 ΨH -1.238290 -0.000271 0.922918 ΨH 1.238290 0.000271 -0.922918 ΨH -1.238290 -0.000271 -0.922918
+S6.8. Initial distorted geometry (in Å) of with cc-pVDZ basis
+with symmetry.
+ΨC 0.670518 0.000007 0.000000 ΨC -0.670518 -0.000007 0.000000 ΨH 1.238283 -0.000164 0.922917 ΨH -1.238283 0.000164 0.922917 ΨH 1.238283 -0.000164 -0.922917 ΨH -1.238283 0.000164 -0.922917